# Evaluating Visual and Cultural Interpretation: The K-Viscuit Benchmark with Human-VLM Collaboration

### v2

@@ Abstract
To create culturally inclusive vision-language models (VLMs), developing a benchmark that tests their ability to address culturally relevant questions is essential. Existing approaches typically rely on human annotators, making the process labor-intensive and creating a cognitive burden in generating diverse questions. To address this, we propose a semi-automated framework for constructing cultural VLM benchmarks, specifically targeting multiple-choice QA. This framework combines human-VLM collaboration, where VLMs generate questions based on guidelines, a small set of annotated examples, and relevant knowledge, followed by a verification process by native speakers. We demonstrate the effectiveness of this framework through the creation of K-Viscuit, a dataset focused on Korean culture. Our experiments on this dataset reveal that open-source models lag behind proprietary ones in understanding Korean culture, highlighting key areas for improvement. We also present a series of further analyses, including human evaluation, augmenting VLMs with external knowledge, and the evaluation beyond multiple-choice QA. [^1] Our dataset is available at https://huggingface.co/datasets/ddehun/k-viscuit.
@@
Yujin Baek11footnotemark: 11     ChaeHun Parkthanks:    Equal contribution1    Jaeseok Kim2     Yu-Jung Heo2    Du-Seong Chang2     Jaegul Choo1 1 KAIST AI     2 KT Corporation {yujinbaek,ddehun,jchoo}@kaist.ac.kr {jaeseok.kim,yj.heo,dschang}@kt.com
@@ 1 Introduction
Recent advances in vision-language models (VLMs) have demonstrated remarkable capabilities in tasks ranging from image captioning to visual question answering. However, these models are predominantly trained on Western-centric datasets (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), leading to significant performance disparities when applied to non-Western contexts (Liu et al., 2021; Yin et al., 2021, 2023; Romero et al., 2024). This cultural bias is particularly problematic as visual interpretation often depends heavily on cultural context, necessitating the development of more culturally aware VLMs.
Several benchmarks have been proposed to evaluate cultural understanding in VLMs (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024; Nayak et al., 2024). These approaches primarily rely on manual question generation, which, while valuable, faces certain practical challenges. The manual process can be time-consuming and resource-intensive when scaling to new cultural contexts. Additionally, as noted in cognitive science research (Ramos, 2020), human annotators may experience cognitive fixation, potentially limiting the diversity of generated questions. These practical considerations motivate the need for more efficient benchmark construction approaches.
Figure 1: Framework Overview.
Inspired by recent successes in human-LLM collaborative data generation (Liu et al., 2022; Bartolo et al., 2022; Kamalloo et al., 2023), we propose a semi-automated framework for constructing cultural VLM benchmarks that enhances both the efficiency and diversity of culture-relevant visual question and answer generation, as shown in Fig. 1. Our framework incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers based on carefully crafted guidelines, a small set of human-annotated examples, and image-specific knowledge. Native speakers then verify these recommended questions to ensure quality and cultural relevance.
Using this framework, we develop K-Viscuit (Korean Visual and Cultural Interpretation Test), a benchmark dataset for Korean culture that can be adapted for other cultural contexts. K-Viscuit features two distinct evaluation types: visual recognition and visual reasoning. Also, the benchmark employs carefully designed multiple-choice questions with highly similar distractors to prevent models from exploiting superficial patterns.
Our evaluation with K-Viscuit reveals a significant performance gap between open-source and proprietary VLMs in understanding Korean culture. We provide insights into the current limitations and potential improvements in VLMs’ cultural understanding capabilities through detailed analyses, which include human evaluation, external knowledge integration, and extended evaluation beyond a multi-choice question-answering setup.
@@ 3.1.1 Concept Categorization
Inspired by recent studies on multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge of various concepts encountered in daily life by Korean natives. While each concept should have some degree of universality, its manifestation often varies across cultures. Following Liu et al. (2021), we reference semantic concepts from the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015) to define our concept list. Our dataset encompasses ten core concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
@@ 3.1.2 Image Selection
Korean native annotators collected web images corresponding to the selected concepts. To ensure diverse representation, we limited each specific object to appearing no more than twice within any single category. Following Liu et al. (2021), we selected only images depicting concepts that could physically exist in everyday life. Annotators were encouraged to source diverse and suitable images from various web resources. Wikimedia Commons [^2] https://commons.wikimedia.org/wiki served as the primary source, and only CC-licensed images were selected.
@@ 3.1.3 Question Generation
Question Type
Based on the selected images, we annotate questions in multiple-choice QA format. To comprehensively evaluate understanding of Korean culture, we categorize questions into two types: visual recognition (Type 1) and reasoning (Type 2). Visual recognition questions assess basic visual information such as object identification, while reasoning questions require fine-grained cultural knowledge or deeper reasoning processes related to the image. For each image, we created one Type 1 question and between one to four Type 2 questions. This categorization offers two key advantages: First, Type 1 questions enable assessment of a model’s basic visual understanding of culturally embedded concepts. Second, Type 2 questions comprehensively evaluate cultural understanding beyond simple object recognition.
@@
AI-assisted Question Annotation
We create questions and their options (one correct answer and three distractors) by leveraging a powerful proprietary VLM (GPT-4-Turbo). For each concept category, human annotators first create exemplar questions and options for at least three images. These manually annotated examples serve as demonstrations for the VLM to generate additional questions and options.
Specifically, the VLM receives: 1) the target image, 2) human-annotated demonstration examples, 3) detailed annotation guidelines, and 4) image-specific knowledge descriptions. We include relevant contextual knowledge for each image to enhance question diversity and relevance, ensuring VLM-generated questions are grounded in real-world understanding. Notably, following Wang et al. (2023), our guidelines emphasize maintaining high similarity among all four multiple-choice options, a principle also reflected in human-annotated examples. All information is provided to the VLM through natural language prompts, with distinct annotation guidelines for visual recognition and reasoning questions. The detailed prompts are presented in Appendix A.
@@
Figure 3: Concept distribution of our dataset.
Model All Food Beverage Game Celeb. Religion Tool Clothes Heritage Arch. Agri. InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Molmo-7B-D (Deitke et al., 2024) 61.04 58.45 71.15 44.23 68.09 44.44 61.33 70.27 56.45 64.22 68.52 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Llama-3.2-11B (Dubey et al., 2024) 68.04 61.27 65.38 50.00 72.34 70.37 72.00 75.68 72.58 69.72 81.48 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19 GPT-4o (OpenAI, 2024) 89.50 88.73 82.69 86.54 95.74 85.19 90.67 91.89 91.94 91.74 87.04
@@ 3.1.4 Human Verification
Our preliminary studies showed that while the VLM often produced factually correct samples, some did not fully align with the cultural nuances we sought to capture. Rather than discarding only incorrect content, our human verification process prioritizes selecting question-option sets that best reflect intended cultural subtleties. Although this leads to setting aside numerous plausible samples, it does not imply unreliability; instead, it refines the dataset to ensure deep cultural resonance. Approved samples require minimal revision before inclusion. This approach emphasizes nuanced cultural alignment and indicates the need for further work to better synchronize VLM outputs with human cultural intentions.
Figure 4: Required Cultural Knowledge.
@@ 3.2 K-Viscuit
Statistics
Table 1 presents detailed statistics of our benchmark dataset. Our dataset comprises 657 total examples (237 Type 1 and 420 Type 2 questions) based on 237 unique images across 10 concept categories. The average word counts are 10.11 and 15.46 for Type 1 and Type 2 questions respectively, with an overall average of 13.53 words. Each question includes four options, totaling 2,628 options with an average length of 1.74 words. Fig. 3 shows the distribution of concept categories in our dataset. We also compare our dataset size with CVQA (Romero et al., 2024), a recently proposed cultural VQA benchmark, in Appendix E.
Required Knowledge to Solve Questions
To characterize our dataset, we analyze the types of knowledge required to solve questions. Following Tong et al. (2024), we used GPT-4 to analyze Type 2 questions. We provided all Type 2 QA pairs and summarized the required knowledge for each question. As shown in Fig. 4, the analysis reveals diverse cultural elements. Detailed categorization instructions are provided in Appendix B.1.
Qualitative Examples
Fig. 2 showcases sample images, questions, and options along with their concept categories and question types.
@@ 4 Experiments
We conduct experiments to evaluate various VLMs on our constructed dataset.
@@ 4.1 Models
The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), Molmo-7B-D (Deitke et al., 2024), Idefics2-8B Laurençon et al. (2024), and Llama-3.2-11B-Vision-Instruct (Dubey et al., 2024). We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), Gemini-1.5-Pro (Reid et al., 2024), and GPT-4o (OpenAI, 2024). All models are evaluated in the conventional multiple-choice setup, where a model is prompted to choose its answer from one of four options. The input text is constructed by concatenating (1) a question, (2) each option with option letters in alphabetical order, and (3) the instruction about output format (i.e., "Answer with the option’s letter from the given choices directly."). We use accuracy as an evaluation metric.
4.2 Results
Table 2 demonstrates the evaluation results of different VLMs on our dataset. We first observe that proprietary models usually show higher accuracies. The GPT-4o and Gemini-1.5-Pro achieve the highest and second-highest scores, respectively. Among the open-sourced models, Llama-3.2-11B and Idefics2-8B usually perform better than other models. The accuracy of models in different question types is shown in Table 3. We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions. Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs. Our dataset enables evaluating and identifying such aspects where models can be improved.
Model Type 1 Type 2 Total
@@ 4.2 Results
Table 3: Results on Different Question Types. Type 1 and Type 2 denote question types in visual recognition and visual reasoning, respectively.
@@ 4.3 Analyses
Figure 5: Human Evaluation. Accuracy comparison between Koreans and non-Koreans on 50 samples of K-Viscuit. The performance of selected models on this subset is also displayed. The average scores for Koreans and non-Koreans are 80.2 and 47.0, respectively.
Human Evaluation on the Benchmark
We conducted a human evaluation to evaluate how well people from different backgrounds perform on K-Viscuit. We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions. This test was administered to 20 Koreans and 14 non-Koreans. As depicted in Fig. 5, the results showed that participants with better knowledge of Korean culture achieved higher accuracy. However, even among Koreans, knowledge gaps exist between individuals and within a single person’s knowledge across different domains, such as history. A Proprietary VLM (i.e., GPT-4-Turbo) shows comparable performance to human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators. Details of the user study are provided in the Appendix B.2.
@@
LLaVA-v1.6-13B 82.8 57.9 57.1
Molmo-7B-D 85.6 65.5 61.0
Idefics2-8b 81.2 69.0 63.6
Llama-3.2-11B 75.2 72.4 68.0
@@
5 Discussion
We discuss several components and future directions to build VLMs grounded on diverse cultures.

### v3

@@ Abstract
To create culturally inclusive vision-language models (VLMs), developing a benchmark that tests their ability to address culturally relevant questions is essential. Existing approaches typically rely on human annotators, making the process labor-intensive and creating a cognitive burden in generating diverse questions. To address this, we propose a semi-automated framework for constructing cultural VLM benchmarks, specifically targeting multiple-choice QA. This framework combines human-VLM collaboration, where VLMs generate questions based on guidelines, a small set of annotated examples, and relevant knowledge, followed by a verification process by native speakers. We demonstrate the effectiveness of this framework through the creation of K-Viscuit, a dataset focused on Korean culture. Our experiments on this dataset reveal that open-source models lag behind proprietary ones in understanding Korean culture, highlighting key areas for improvement. We also present a series of further analyses, including human evaluation, augmenting VLMs with external knowledge, and the evaluation beyond multiple-choice QA. Our dataset is available at https://huggingface.co/datasets/ddehun/k-viscuit.
@@
ChaeHun Park11footnotemark: 11     Yujin Baekthanks:    Equal contribution1    Jaeseok Kim2     Yu-Jung Heo2    Du-Seong Chang3     Jaegul Choo1 1 KAIST AI     2 KT Corporation     3 Sogang Univ. {ddehun,yujinbaek,jchoo}@kaist.ac.kr {jaeseok.kim,yj.heo}@kt.com     duseong.chang@gmail.com
@@ 1 Introduction
Recent advances in vision-language models (VLMs) have demonstrated remarkable capabilities in tasks ranging from image captioning to visual question answering. However, these models are predominantly trained on Western-centric datasets (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), leading to substantial performance disparities when applied to non-Western contexts (Liu et al., 2021; Yin et al., 2021, 2023; Romero et al., 2024). This cultural bias is particularly problematic as visual interpretation often depends heavily on cultural context, necessitating the development of more culturally aware VLMs.
Several benchmarks have been proposed to evaluate cultural understanding in VLMs (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024; Nayak et al., 2024; Bhatia et al., 2024). These approaches primarily rely on manual question generation, which, while valuable, faces certain practical challenges. The manual process can be time-consuming and resource-intensive when scaling to new cultural contexts. Additionally, as noted in cognitive science research (Ramos, 2020), humans may experience cognitive fixation, potentially limiting the diversity of generated questions. These practical considerations motivate the need for more efficient benchmark construction approaches.
Inspired by recent successes in human-LLM collaborative data generation (Liu et al., 2022; Bartolo et al., 2022; Kamalloo et al., 2023), we propose a semi-automated framework for constructing cultural VLM benchmarks that enhances both the efficiency and diversity of culture-relevant visual question and answer generation (see Fig. 2). Our framework incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers based on carefully crafted guidelines, a small set of human-annotated examples, and image-specific knowledge. Native speakers then verify these recommended questions to ensure quality and cultural relevance.
Using this framework, we develop K-Viscuit (Korean Visual and Cultural Interpretation Test), a benchmark dataset tailored to Korean culture. Note that the framework is adaptable and can be similarly applied to other cultures. K-Viscuit features two distinct evaluation types: visual recognition and visual reasoning. Also, the benchmark employs carefully designed multiple-choice questions with highly similar distractors to prevent models from exploiting superficial patterns.
Our evaluation with K-Viscuit reveals a notable performance gap between open-source and proprietary VLMs in understanding Korean culture. We provide insights into the current limitations and potential improvements in VLMs’ cultural understanding capabilities through detailed analyses, which include human evaluation, external knowledge integration, and extended evaluation beyond a multi-choice question-answering setup.
@@
Figure 2: Framework Overview. We begin with image selection, where concept-relevant images (e.g., the traditional Korean game Yutnori) are chosen and briefly described. In the QA generation stage, a VLM is guided by annotation guidelines and demonstration samples to draft question–option pairs. Finally, human verification ensures annotator review, correction, and filtering of the generated samples.
@@ 3.1.1 Concept Categorization
Inspired by recent studies on multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge of various concepts encountered in daily life by Korean natives. While each concept should have some degree of universality, its manifestation often varies across cultures. Following Liu et al. (2021), we reference semantic concepts from the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015) to define our concept list. Our dataset encompasses ten core concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture. Further details are in Appendix B.1.
@@ 3.1.2 Image Selection
Korean native annotators collected web images corresponding to the selected concepts. To ensure diverse representation, we limited each specific object to appearing no more than twice within any single category. Following Liu et al. (2021), we selected only images depicting concepts that could physically exist in everyday life. Annotators were encouraged to source diverse and suitable images from various web resources. [^1] Although framed as image selection, this step also served as a concept-level filtering process. For instance, we excluded abstract or symbolic visuals (e.g., fireworks emojis for “celebration”) that lacked concrete visual referents. Wikimedia Commons [^2] https://commons.wikimedia.org/wiki served as the primary source, and only CC-licensed images were selected.
@@ 3.1.3 Question Generation
Question Type Based on the selected images, we annotate questions in a multiple-choice QA format. To comprehensively evaluate the understanding of Korean culture, we categorize questions into two types: Visual Recognition (Type 1) and Cultural Knowledge Application (Type 2). Type 1 questions assess basic visual information, such as object identification. In contrast, Type 2 questions require a deeper understanding of cultural context, reasoning, or multi-step inference based on the image. For each image, we created one Type 1 question and between one and four Type 2 questions. This categorization provides two key advantages: First, Type 1 questions assess a model’s ability to recognize culturally embedded visual elements. Second, Type 2 questions evaluate cultural understanding beyond simple object recognition.
@@
Figure 3: Concept distribution of our dataset.
AI-assisted Question Annotation We create questions and their options (one correct answer and three distractors) by leveraging a powerful proprietary VLM (GPT-4-Turbo). For each concept category, human annotators first create exemplar questions and options for at least three images. These manually annotated examples serve as demonstrations for the VLM to generate additional questions and options.
Specifically, the VLM receives: 1) the target image, 2) human-annotated demonstration examples, 3) detailed annotation guidelines, and 4) image-specific knowledge descriptions. We include relevant contextual knowledge for each image to enhance question diversity and relevance, ensuring VLM-generated questions are grounded in real-world understanding. Notably, following Wang et al. (2023), our guidelines emphasize maintaining high similarity among all four multiple-choice options, a principle also reflected in human-annotated examples. All information is provided to the VLM through natural language prompts, with distinct annotation guidelines for visual recognition and application questions. The detailed prompts are presented in Appendix B.
Figure 4: Required Cultural Knowledge.
@@
Model All Food Beverage Game Celeb. Religion Tool Clothes Heritage Arch. Agri. InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024a) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024a) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Molmo-7B-D (Deitke et al., 2024) 61.04 58.45 71.15 44.23 68.09 44.44 61.33 70.27 56.45 64.22 68.52 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Llama-3.2-11B (Dubey et al., 2024) 68.04 61.27 65.38 50.00 72.34 70.37 72.00 75.68 72.58 69.72 81.48 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19 GPT-4o (OpenAI, 2024) 89.50 88.73 82.69 86.54 95.74 85.19 90.67 91.89 91.94 91.74 87.04
@@ 3.1.4 Human Verification
In this step, we evaluate whether the generated QA pairs adhere not only to the detailed prompt guidelines (Tables 10 and 11), but also to the intended cultural context. Our preliminary studies revealed that although VLMs often produced factually accurate outputs, some failed to capture the nuanced cultural elements we aimed to represent. Rather than discarding only incorrect content, our human verification process prioritizes selecting question-option sets that best reflect intended cultural subtleties. Although this leads to setting aside numerous plausible samples, it does not imply unreliability; instead, it refines the dataset to ensure deep cultural resonance. In particular, for Type 2 questions, we removed those that primarily tested simple visual recognition rather than cultural knowledge application. Approved samples require minimal revision before inclusion, emphasizing nuanced cultural alignment and highlighting the need for further work to better synchronize VLM outputs with human cultural intentions.
@@ 3.2 K-Viscuit
Statistics Table 1 presents detailed statistics of our benchmark dataset. Our dataset comprises 657 total examples (237 Type 1 and 420 Type 2 questions) based on 237 unique images across 10 concept categories. The average word counts are 10.11 and 15.46 for Type 1 and Type 2 questions respectively, with an overall average of 13.53 words. Each question includes four options, totaling 2,628 options with an average length of 1.74 words. Fig. 3 shows the distribution of concept categories in our dataset. We also compare our dataset size with CVQA (Romero et al., 2024), a recently proposed cultural VQA benchmark, in Appendix A.2.
Required Knowledge to Solve Questions To characterize our dataset, we analyze the types of knowledge required to solve questions. Following Tong et al. (2024), we used GPT-4 to analyze Type 2 questions. We provided all Type 2 QA pairs and summarized the required knowledge for each question. As shown in Fig. 4, the analysis reveals diverse cultural elements. Detailed categorization instructions are provided in Appendix C.1.
Qualitative Examples Fig. 1 showcases sample images, questions, and options along with their concept categories and question types.
@@ 4 Experiments
This section presents our experiments, including VLM evaluation, human benchmarking, and the impact of question language on accuracy.
@@ 4.1 Models
The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024a), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), Molmo-7B-D (Deitke et al., 2024), Idefics2-8B Laurençon et al. (2024), and Llama-3.2-11B-Vision-Instruct (Dubey et al., 2024). We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), Gemini-1.5-Pro (Reid et al., 2024), and GPT-4o (OpenAI, 2024). All models are evaluated in the conventional multiple-choice setup, where a model is prompted to choose its answer from one of four options. The input text is constructed by concatenating (1) a question, (2) each option with option letters in alphabetical order, and (3) the instruction about output format (i.e., "Answer with the option’s letter from the given choices directly."). We use accuracy as an evaluation metric.
4.2 Results and Analyses
Main Results
Table 2 demonstrates the evaluation results of different VLMs on our dataset. We first observe that proprietary models usually show higher accuracies. The GPT-4o and Gemini-1.5-Pro achieve the highest and second-highest scores, respectively. Among the open-sourced models, Llama-3.2-11B and Idefics2-8B usually perform better than other models. The accuracy of models in different question types is shown in Table 3. We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions. Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs. [^3] Further analysis of this trend is provided in A.3. Our dataset enables evaluating and identifying such aspects where models can be improved.
Model Type 1 Type 2 All
@@ 4.2 Results and Analyses
Table 3: Results on Different Question Types. Type 1 and Type 2 denote question types in visual recognition and cultural knowledge application, respectively.
Figure 5: Human Evaluation. Accuracy comparison between Koreans and non-Koreans on 50 samples of K-Viscuit. The performance of selected models on this subset is also displayed. The average scores were 80.2 (Korean participants) and 47.0 (non-Korean participants), with standard deviations of 2.69 and 5.95.
Human Evaluation on the Benchmark
We conducted a human evaluation to evaluate how well people from different backgrounds perform on K-Viscuit. We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions. This test was administered to 20 Koreans and 14 non-Koreans. As depicted in Fig. 5, the results showed that participants with better knowledge of Korean culture achieved higher accuracy. However, even among Koreans, knowledge gaps exist between individuals and within a single person’s knowledge across different domains, such as history. A Proprietary VLM (i.e., GPT-4-Turbo) shows comparable performance to human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators. Details of the user study are provided in the Appendix C.2.
@@
LLaVA-v1.6-13B 82.8 57.93 57.08
Molmo-7B-D 85.6 65.52 61.04
Idefics2-8b 81.2 68.97 63.62
Llama-3.2-11B 75.2 72.41 68.04
@@
Figure 7: Visual Dependency Analysis. Accuracy comparison of selected models when using actual images versus noise images. The numbers indicate absolute and relative accuracy drops when replacing images with noise.
Evaluating the Visual Dependency of K-Viscuit Questions
To assess how much the K-Viscuit dataset relies on visual information, we conducted an experiment measuring the impact of removing meaningful image content. Specifically, models are provided with randomly generated images with pixel values sampled from a Gaussian distribution instead of actual images. These images are of the same size as the originals but contain no meaningful visual structure. By comparing model performance in this setting to standard image-based evaluation, we can determine the extent to which K-Viscuit questions require visual understanding rather than relying solely on textual cues.
Fig. 7 presents the accuracy of selected models under two conditions: with actual images (Image + Text) and with noise images (Text Only). Across all models, a substantial drop in accuracy is observed when images are replaced with noise, confirming that K-Viscuit questions require meaningful visual information. The performance degradation varies by model, with Llama-3.2-11B showing the largest accuracy drop, while Molmo-7B-D is comparatively less affected. Additionally, both Type 1 and Type 2 questions exhibit large performance declines in the absence of images, suggesting that visual information is crucial across different question types. These results highlight the varying degrees of visual dependence across models and reinforce the necessity of strong visual reasoning capabilities for solving K-Viscuit questions.
@@
5 Future Directions
This section explores future directions for improving model evaluation (§5.1) and enhancing model performance with external knowledge (§5.2).